OpenStudy (anonymous):

(x + 1)2 is equal to which of the following? a.)(x + 1)(x - 1) b.)(x + 1)(x + 1) c.)x2 + 1 d.) x2 - 1

OpenStudy (anonymous):

no not really cans you teach me ^.^

OpenStudy (anonymous):

LOL you don't even have to foil here

OpenStudy (anonymous):

so when you have something raised to the 2nd power, it just means that you are multiplying it by itself just once for example\[a^{2} = a * a\]

OpenStudy (anonymous):

so\[(x+1)^{2} = (x+1)(x+1)\]